# Extreme events and power-law distributions from nonlinear quantum dissipation

### v1

@@ Abstract
where $\nu=(\gamma_{1}-\kappa_{1})/4\gamma_{2}$ is the power-law exponent. Equation (5) is valid only for $\nu>2$ [30], and plotted in Fig. 2(a). The predicted divergence at $\nu=2$ is consistent with the power-law tail (3). For $\nu<2$, the mean excitation number $\sim\exp[(2-\nu)4\gamma_{2}t]$ blows up exponentially in time. The second moment can also be computed in the steady state as
@@
for $\nu>3$, and blows up exponentially $\sim\exp[(3-\nu)8\gamma_{2}t]$ for $\nu<3$. Again, the divergence at $\nu=3$ is consistent with the power-law tail. All higher moments can be computed recursively [30].
@@
This has a minimum value of $g^{(2)}(0)=8$ at $\nu=6$, which reflects the fact that the steady state is not at thermal equilibrium. $g^{(2)}(0)$ diverges at $\nu=3$ due to the strong fluctuations in the bosonic field. For $\nu\gg 1$, $g^{(2)}(0)\sim(\nu+7)/2$ grows linearly with $\nu$ [see Fig. 2(c)]. This regime could be accessible in practical implementations where $\{\gamma_{2},\kappa_{2}\}\ll\gamma_{1}$. Thus, the linear behavior of $g^{(2)}(0)$ can potentially serve as an experimental signature of the power-law tail. We remark that the $k$-boson correlation function analogous to Eq. (7) scales as $g^{(k)}(0)\equiv\braket{\hat{a}^{{\dagger}k}\hat{a}^{k}}/\braket{\hat{a}^{{\dagger}}\hat{a}}^{k}\sim\nu^{\left\lfloor{k/2}\right\rfloor}$ for large $\nu$ [30].
The behavior of $g^{(2)}(0)$ for $\nu\leq 3$ is less straightforward to analyze. Obviously, the correlation function is infinite in the regime $2<\nu\leq 3$, since $\braket{\hat{a}^{{\dagger}}\hat{a}}$ is finite, while $\braket{\hat{a}^{{\dagger}2}\hat{a}^{2}}$ diverges. However, for $\nu\leq 2$, both $\braket{\hat{a}^{{\dagger}2}\hat{a}^{2}}$ and $\braket{\hat{a}^{{\dagger}}\hat{a}}$ diverge at the critical point. We can gain insights about the physics for $\nu<2$ using a simple scaling theory [30]: Suppose $\rho_{n,n}\propto n^{-\nu}$ has a sharp cutoff at $n_{c}=1/\delta$. A straightforward calculation predicts the scaling behavior $\braket{(\hat{a}^{{\dagger}}\hat{a})^{k}}\sim 1/\delta^{k}$ for $\nu\in(0,1)$ and $\sim 1/\delta^{k-\nu+1}$ for $\nu\in(1,k+1)$, with $k\in\mathbb{Z}^{+}$. Consequently, $g^{(2)}(0)$ diverges as $1/\delta^{\nu-1}$ for $\nu\in(1,2)$. Replacing the sharp cutoff with the exponential cutoff in Eq. (3) only adds a finite correction and does not affect the scaling of the divergences. The predicted divergences from the scaling theory agree with numerical simulations near the critical point, as shown in Fig. 2(b). Note that the observed behavior in Fig. 2(b) cannot be extrapolated to the critical point $\delta=0$, since $g^{(2)}(0)\sim\exp(8\gamma_{2}t)$ blows up exponentially in time for all $\nu<2$. Mathematically, this corresponds to a singular perturbation. In the regime $\nu\ll\delta$, the steady state is close to thermal equilibrium and thus $g^{(2)}(0)\approx 2$, consistent with the numerical observation.
The power-law behavior can also manifest in the quantum phase space. Introducing the dimensionless quadrature operators $\hat{q}=(\hat{a}+\hat{a}^{{\dagger}})/\sqrt{2}$ and $\hat{p}=-i(\hat{a}-\hat{a}^{{\dagger}})/\sqrt{2}$, with the canonical commutation relation $[\hat{q},\hat{p}]=i$, we show analytically that the Wigner quasiprobability distribution in the steady state has the asymptotic behavior [30]
@@
with the same power-law exponent $\nu$ as the number distribution. This can be observed, for example, by performing tomography on the bosonic mode. If the system represents a quantum particle confined in a one-dimensional harmonic potential, the position and momentum distribution of the particle exhibits the tail behavior $\sim|q|^{1-2\nu}$ and $\sim|p|^{1-2\nu}$ respectively, from Eq. (8).
@@
To summarize, the power-law distribution due to quantum fluctuations gives rise to extreme events which can be orders of magnitude above the typical measured values. As shown above, such effects can be observed through various physical quantities that have a scaling behavior dependent on the power-law exponent $\nu$.
General class of nonlinear dissipative dynamics.— Now, we argue that the power-law behavior is not specific to the $M$-boson model of Eq. (1), but is instead a generic feature of quantum systems with strong nonlinear dissipation. To this end, we consider a general class of quantum systems that obey the equation of motion
$\begin{split}\frac{d\braket{\hat{q}}}{dt}&=\braket{\hat{p}},\\ \frac{d\braket{\hat{p}}}{dt}&=\left(k_{0}\braket{\hat{p}}-k_{2}\braket{:{\hat{q}^{2}\hat{p}}:}\right)-\left(k_{1}\braket{\hat{q}}+k_{3}\braket{:\hat{q}^{3}:}\right),\end{split}$ (9)
@@
reminiscent of the Ehrenfest theorem. Here, $\hat{q}$ and $\hat{p}$ are the quadrature operators defined above, while $\{k_{0},k_{1},k_{2},k_{3}\}$ are arbitrary non-negative parameters. The notation $\braket{:\hat{O}:}$ indicates normal ordering, after expressing the operator $\hat{O}$ in terms of $\hat{a}$ and $\hat{a}^{{\dagger}}$. The choice of operator ordering here is simply for convenience, and does not affect the discussion. Eq. (9) is the quantum analog of the cubic Liénard system. Physically, the parameters $k_{2}$ and $k_{3}$ represent the nonlinear damping and frequency response (i.e., anharmonicity) of the system, respectively.
@@
Figure 3: Magnitudes of density matrix element $|\rho_{n,n^{\prime}}|$ against the number of excitation quanta $n$ for the quantum Liénard system (9) in the steady state. Curves plotted are for even $n^{\prime}-n\in[0,16]$. Matrix elements for odd $n^{\prime}-n$ are identically zero. The black dashed line shows the power-law scaling $|\rho_{n,n^{\prime}}|\propto n^{-1.6}$ obtained from numerical fitting the number distribution $\rho_{n,n}$. (Inset) Wigner quasiprobability distribution $W(q,p)$ of the steady state. $W(q,p)$ is mostly supported near the phase space origin, with a heavy tail that decays slowly. Parameters: $(k_{0},k_{1},k_{2},k_{3})=(0.169,0.12,1,0.035)$, with a Hilbert space truncation of $N_{T}=10^{3}$ in the numerical simulation for numerical convergence. Similar behavior can be observed for other parameters.
@@
We seek to ‘quantize’ the Liénard system by constructing a Lindblad master equation that generates the quantum dynamics (9). A general construction for any polynomial dynamical system with a two-dimensional phase space is detailed in Refs. [21, 22]. The resulting master equation is rather complicated, and is included in the SM [30]. The key qualitative difference with the $M$-boson model (1) is that the dynamical system (9) does not possess continuous rotation symmetry. Therefore, the steady state (if it exists) will contain coherences in the excitation number basis. In the regime of weak nonlinear dissipation $k_{2}\ll k_{1}$, the master equation can be described by a quantum Stuart-Landau model with a generalized squeezing drive [33, 34, 30], which has an exponentially vanishing number distribution. Thus, strong nonlinear dissipation is needed to observe the power-law tail.
@@
We provide numerical evidence for the emergence of power-law tails in the steady state. Unlike the $M$-boson model (1), the power law in the number distribution $\rho_{n,n}$ can be observed without fine-tuning of the parameters. We demonstrate this by simulating the master equation with parameters $k_{0},k_{1}$ and $k_{3}$ randomly sampled from $[0,1]$ (in units of $k_{2}$). An explicit example is shown in Fig. 3. Interestingly, the power law does not only emerge in the probabilities $\rho_{n,n}$, like for the $M$-boson model, but also in the off-diagonal elements (coherences) of the density matrix. In the example shown, we observe the scaling $|\rho_{n,n^{\prime}}|\sim n^{-1.6}$ for $n^{\prime}-n$ up to $16$. This is in the regime of infinite mean excitation number, akin to $\nu<2$ in the $M$-boson model. The physical implication is that extreme events can be generically observed in quantum systems with strong nonlinear dissipation, due to the amplified quantum fluctuations.
@@
Discussion.— We propose the emergence of power-law distributions and extreme events as a novel universal feature of nonlinear quantum dissipation. This can be understood from the fact that the quantum noise associated with nonlinear dissipation is typically multiplicative in nature. That is, the fluctuations are coupled to the quantum state and become amplified at higher excitation. The competition between dissipation and noise amplification results in the power-law behavior. We introduce the prototypical $M$-boson model which is analytically proven to produce power-law energy distributions at the critical point. The ubiquity of this phenomenon is further supported via numerical simulations on a general class of nonlinear dynamical systems which are quantum analogs of Liénard systems. We show that power laws can also emerge in the (off-diagonal) coherences of the density matrix, which exhibit the same scaling as the (diagonal) probability distribution.
The power-law behavior can manifest in various observable physical signatures, such as extreme events with divergent mean excitation number, boson superbunching with a divergent $g^{(2)}(0)$ correlation function, and power-law tails in both position and momentum space. This phenomenon opens up potential avenues for exploring extreme regimes of light-matter interactions [35, 36, 37], and can be a useful resource for sensing applications like ghost imaging [38, 39]. It would also be interesting as a future theoretical study to connect the results of this work with known mechanisms of the power law [1], such as scale invariance, universality and self-organized criticality.
@@
Acknowledgements.— The author would like to thank Andy Chia, Tobias Haug, Leong-Chuan Kwek, Changsuk Noh, Jingu Pang, John Preskill, Gil Refael and Yu Tong for insightful discussions. The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.
Supplemental Material
We provide additional technical details supporting the claims in the main text.
I Steady state number distribution for the $M$-boson model
@@
This describes a quantum harmonic oscillator subject to incoherent processes such as damping (governed by the $\gamma_{m}$ terms) and pumping (governed by the $\kappa_{m}$ terms). On top of the harmonic oscillator Hamiltonian $H_{0}=\hbar\omega_{0}\hat{a}^{{\dagger}}\hat{a}$, one can also consider a generalized model by adding any Hamiltonian which is diagonal in the number (Fock) basis {|$\{\ket{n}\}$. This includes, for example, a Kerr nonlinearity $H_{\text{Kerr}}=K\hat{a}^{{\dagger}2}\hat{a}^{2}$ which controls the anharmonicity of the oscillator. Such terms do not contribute to the dynamics of the populations (diagonal elements of $\rho$ in the number basis),
@@
II Analytical results for $M=2$ at the critical point
Here, we present additional analytical results for the $M$-boson model.
II.1 Moments of the number distribution

### v2

@@ 3 Physical signatures of the power-law tail
where $\nu=(\gamma_{1}-\kappa_{1})/4\gamma_{2}$ is the power-law exponent, see Appendix B.1. Equation (5) is valid only for $\nu>2$, and is plotted in Fig. 2(a). The predicted divergence at $\nu=2$ is consistent with the power-law tail (3). For $\nu<2$, the mean excitation number $\sim\exp[(2-\nu)4\gamma_{2}t]$ blows up exponentially in time. The second moment can also be computed in the steady state as
@@
for $\nu>3$, and blows up exponentially $\sim\exp[(3-\nu)8\gamma_{2}t]$ for $\nu<3$. Again, the divergence at $\nu=3$ is consistent with the power-law tail. All higher moments can be computed recursively, as derived in Appendix B.1.
@@
This has a minimum value of $g^{(2)}(0)=8$ at $\nu=6$, which reflects the fact that the steady state is not at thermal equilibrium. $g^{(2)}(0)$ diverges at $\nu=3$ due to the strong fluctuations in the bosonic field. For $\nu\gg 1$, $g^{(2)}(0)\sim(\nu+7)/2$ grows linearly with $\nu$ [see Fig. 2(c)]. This regime could be accessible in practical implementations where $\{\gamma_{2},\kappa_{2}\}\ll\gamma_{1}$. Thus, the linear behavior of $g^{(2)}(0)$ can potentially serve as an experimental signature of the power-law tail. We remark that the $k$-boson correlation function analogous to Eq. (7) scales as $g^{(k)}(0)\equiv\braket{\hat{a}^{{\dagger}k}\hat{a}^{k}}/\braket{\hat{a}^{{\dagger}}\hat{a}}^{k}\sim\nu^{\left\lfloor{k/2}\right\rfloor}$ for large $\nu$, see Appendix B.2.
The behavior of $g^{(2)}(0)$ for $\nu\leq 3$ is less straightforward to analyze. Obviously, the correlation function is infinite in the regime $2<\nu\leq 3$, since $\braket{\hat{a}^{{\dagger}}\hat{a}}$ is finite, while $\braket{\hat{a}^{{\dagger}2}\hat{a}^{2}}$ diverges. However, for $\nu\leq 2$, both $\braket{\hat{a}^{{\dagger}2}\hat{a}^{2}}$ and $\braket{\hat{a}^{{\dagger}}\hat{a}}$ diverge at the critical point. We can gain insights about the physics for $\nu<2$ using a simple scaling theory (see Appendix C for details): Suppose $\rho_{n,n}\propto n^{-\nu}$ has a sharp cutoff at $n_{c}=1/\delta$. A straightforward calculation predicts the scaling behavior $\braket{(\hat{a}^{{\dagger}}\hat{a})^{k}}\sim 1/\delta^{k}$ for $\nu\in(0,1)$ and $\sim 1/\delta^{k-\nu+1}$ for $\nu\in(1,k+1)$, with $k\in\mathbb{Z}^{+}$. Consequently, $g^{(2)}(0)$ diverges as $1/\delta^{\nu-1}$ for $\nu\in(1,2)$. Replacing the sharp cutoff with the exponential cutoff in Eq. (3) only adds a finite correction and does not affect the scaling of the divergences. The predicted divergences from the scaling theory agree with numerical simulations near the critical point, as shown in Fig. 2(b). Note that the observed behavior in Fig. 2(b) cannot be extrapolated to the critical point $\delta=0$, since $g^{(2)}(0)\sim\exp(8\gamma_{2}t)$ blows up exponentially in time for all $\nu<2$. Mathematically, this corresponds to a singular perturbation. In the regime $\nu\ll\delta$, the steady state is close to thermal equilibrium and thus $g^{(2)}(0)\approx 2$, consistent with the numerical observation.
The power-law behavior can also manifest in the quantum phase space. Introducing the dimensionless quadrature operators $\hat{q}=(\hat{a}+\hat{a}^{{\dagger}})/\sqrt{2}$ and $\hat{p}=-i(\hat{a}-\hat{a}^{{\dagger}})/\sqrt{2}$, with the canonical commutation relation $\hat{q},\hat{p}]=i$, we show analytically in Appendix B.3 that the Wigner quasiprobability distribution in the steady state has the asymptotic behavior
@@
with the same power-law exponent $\nu$ as the number distribution. This can be observed, for example, by performing tomography on the bosonic mode. If the system represents a quantum particle confined in a one-dimensional harmonic potential, the position and momentum distributions of the particle exhibit the tail behavior $\sim|q|^{1-2\nu}$ and $\sim|p|^{1-2\nu}$ respectively, from Eq. (8).
@@
To summarize, the power-law distribution due to quantum fluctuations gives rise to extreme events which can be orders of magnitude above the typical measured values. As shown above, such effects can be observed through various physical quantities that have a scaling behavior dependent on the power-law exponent $\nu$. We emphasize that the behavior of the various physical quantities shown in Fig. 2 is not intrinsic to the $M$-boson model. A similar behavior is expected for any quantum system exhibiting a power-law tail $\rho_{n,n}\sim n^{-\nu}$, although the value of $\nu$ may be model-dependent.
4 Power-law behavior in general nonlinear dynamics
The $M$-boson model serves as a toy model to gain insights about the physical origin of the power-law tail. However, one might argue that the phenomenon is not truly quantum mechanical, since Eq. (2) is a classical master equation, and multiplicative noise is well-known to produce power-law distributions in classical systems [43, 44]. Thus, the more interesting question is: Are there dissipative quantum systems that exhibit power-law tails in the steady state not only in the populations, but also in the coherences (i.e., off-diagonal elements of $\rho$)? If so, this would indicate that the intrinsically quantum nature of the noise plays an important role.
To this end, let us consider a general classical dynamical system of the form
$\frac{d^{2}q}{dt^{2}}=F(q,\dot{q}),$ (9)
where $F$ is an arbitrary function describing the force acting on a classical particle. We then construct its quantization, i.e., a Lindblad master equation that yields the generalized Ehrenfest theorem
$\frac{d\braket{\hat{q}}}{dt}=\braket{\hat{p}},\quad\frac{d\braket{\hat{p}}}{dt}=\braket{F(\hat{q},\hat{p})}$, (10)
with some suitably defined operator ordering. Here, $\hat{q}$ and $\hat{p}$ are the quadrature operators previously defined. For the evolution to be physically valid, we impose the constraint that the master equation must generate dynamical maps that are completely positive and trace preserving (CPTP) [45, 15]. This general quantization problem was recently solved in Ref. [19], for arbitrary bivariate polynomials $F(q,p)$. A rich subclass of dynamical systems is the cubic Liénard system [20], with the quantum equation of motion given by
$\begin{split}\frac{d\braket{\hat{q}}}{dt}&=\braket{\hat{p}},\\ \frac{d\braket{\hat{p}}}{dt}&=k_{0}\braket{\hat{p}}-k_{2}\braket{:{\hat{q}^{2}\hat{p}}:}-k_{1}\braket{\hat{q}}-k_{3}\braket{:\hat{q}^{3}:}.\end{split}$ (11)
@@ 4 Power-law behavior in general nonlinear dynamics
The notation $\braket{:\hat{O}:}$ indicates normal ordering, after expressing the operator $\hat{O}$ in terms of $\hat{a}$ and $\hat{a}^{{\dagger}}$. The choice of operator ordering here is simply for convenience, and does not affect the discussion [19]. The cubic Liénard system encompasses paradigmatic nonlinear systems such as the van der Pol, Rayleigh, and Duffing oscillators [20]. Physically, the parameters $k_{2}$ and $k_{3}$ represent the nonlinear damping and nonlinear frequency response (i.e., anharmonicity) of the system, respectively. In the absence of the nonlinearity, i.e., $k_{2}=k_{3}$ = 0, the cubic Liénard system reduces to the damped harmonic oscillator.
@@
Figure 3: Power-law tails in the quantum Liénard system. Magnitudes of steady-state density matrix element $|\rho_{n,n^{\prime}}|$ against the number of excitation quanta $n$ for the quantum Liénard system (11) in the limit-cycle regime. Curves plotted are for even $n^{\prime}-n\in[0,16]$. Matrix elements for odd $n^{\prime}-n$ are identically zero. The black dashed line shows the power-law scaling $|\rho_{n,n^{\prime}}|\propto n^{-1.6}$ obtained from fitting the number distribution $\rho_{n,n}$. (Inset) Wigner quasiprobability distribution $W(q,p)$ of the steady state. $W(q,p)$ is mostly supported near the phase space origin, with a heavy tail that decays slowly. Parameters: $(k_{0},k_{1},k_{2},k_{3})=(0.169,0.12,1,0.035)$, with a Hilbert space truncation of $N_{T}=10^{3}$ in the simulation for numerical convergence. Similar behavior can be observed for other parameters.
Using the quantization method developed in Ref. [19], we obtain the Lindblad master equation corresponding to Eq. (11),
$\dot{\rho}=$ $-i[\hat{H},\rho]+k_{0}\mathcal{D}[\hat{a}^{{\dagger}}]\rho+\frac{k_{2}}{2}\mathcal{D}[\hat{a}^{{\dagger}}\hat{a}-\hat{a}^{{\dagger}2}/2]\rho$ (12)
$+\frac{3k_{2}}{8}\mathcal{D}[\hat{a}^{2}]\rho,$
with the Hamiltonian
$\hat{H}=\hat{H}_{0}+\hat{H}_{1},$ (13)
and
$\hat{H}_{1}$ $=\frac{i}{4}\left[k_{0}-i(k_{1}-1)\right]\hat{a}^{2}-\frac{i}{4}\left(\frac{k_{2}}{2}+ik_{3}\right)\hat{a}^{{\dagger}}\hat{a}^{3}$ (15)
$-\frac{i}{16}(k_{2}+ik_{3})\hat{a}^{4}+\text{h.c.},$
where h.c. denotes the Hermitian conjugate of all the preceding terms.
Although the quantum models are constructed phenomenologically, they can potentially be realized using quantum dissipation engineering (also known as reservoir engineering) [46, 47, 48], which has been widely successful for implementing nonlinear dissipation [37, 38, 49, 50, 39, 40, 51, 35]. The key idea is to engineer a nonlinear Hamiltonian interaction between the system and a highly dissipative ancillary system, which mediates an effective nonlinear dissipation in the system [52]. A wide range of nonlinear dissipation terms can be implemented this way in circuit QED, using asymmetrically threaded SQUIDs (ATS) or similar architectures [38, 53, 54, 55]. Alternatively, one can use continuous measurement and feedback to realize the nonlinear dissipation [56], which was proposed to implement a variant of Eq. (67) [21].
@@
A key qualitative difference with the $M$-boson model (1) is that the dynamical system (11) does not possess continuous rotational symmetry. Therefore, the steady state (if it exists) will contain coherences in the excitation number basis. In the regime of weak nonlinear dissipation $k_{2}\ll k_{1}$, the master equation can be described by a quantum Stuart-Landau model with a generalized squeezing drive [57, 23], which has an exponentially vanishing number distribution (see Appendix D). Thus, strong nonlinear dissipation is needed to observe the power-law tail.
@@
In the classical case, for parameters $k_{i}>0$, $i=0,1,2$ and $k_{3}\geq 0$, Liénard’s theorem guarantees a unique stable limit cycle surrounding the origin [20]. Correspondingly, quantum limit cycles also exist in the quantum system [21, 22, 23]. We provide numerical evidence for the emergence of power-law tails in the quantum steady state. Unlike the $M$-boson model (1), the power law in the number distribution $\rho_{n,n}$ can be observed without fine-tuning of the parameters. We demonstrate this by simulating the master equation with parameters $k_{0},k_{1}$ and $k_{3}$ randomly sampled from $[0,1]$ (in units of $k_{2}$). An explicit example is shown in Fig. 3, with additional numerical results presented in Appendix D.3. Interestingly, the power law emerges not only in the probabilities $\rho_{n,n}$, like for the $M$-boson model, but also in the off-diagonal elements (coherences) of the density matrix. In the example shown, we observe the scaling $|\rho_{n,n^{\prime}}|\sim n^{-1.6}$ for $n^{\prime}-n$ up to $16$. This is in the regime of infinite-mean excitation number, akin to $\nu<2$ in the $M$-boson model. This example suggests that the quantum effects of the fluctuations are relevant. A qualitatively similar behavior is observed for $k_{0}=k_{1}=k_{3}=0$ and $k_{2}=1$, see Appendix D.3. This supports the intuition that the nonlinear dissipation term governed by $k_{2}$ is the key mechanism responsible for the emergence of power-law behavior. While beyond the scope of this work, a systematic analysis of the quantum fluctuations in such models would be valuable. This can potentially provide an analytical explanation for the observed power laws in our numerical simulations, and elucidate how the power-law exponent depends on the parameters.
5 Discussion
@@
In this work, we propose the emergence of power-law distributions and extreme events as a novel feature of nonlinear quantum dissipation. This can be understood from the fact that the quantum noise associated with nonlinear dissipation is typically multiplicative in nature. That is, the fluctuations are coupled to the quantum state and become amplified at higher excitation. The competition between dissipation and noise amplification results in the power-law behavior. We introduce the prototypical $M$-boson model which is analytically proven to produce power-law energy distributions at the critical point. The ubiquity of this phenomenon is further supported via numerical simulations on a general class of nonlinear dynamical systems which are quantum analogs of Liénard systems. We show that power laws can also emerge in the (off-diagonal) coherences of the density matrix, which exhibit the same scaling as the (diagonal) probability distribution.
Conceptually, the emergence of power-law tails also calls into question how classical dynamics arise from quantum dissipative systems. In standard examples such as the damped harmonic oscillator and the quantum Stuart-Landau model [33, 36], the quantum fluctuations are relatively insignificant at large amplitudes (which can be formalized using the system-size expansion [16]), giving a well-defined classical limit. On the contrary, the presence of power-law tails originating from multiplicative quantum noise implies that the quantum fluctuations are not negligible in such cases, even at large amplitudes. Reconciling nonlinear dynamical systems with open quantum systems thus remains an interesting open problem. It would also be interesting as a future theoretical study to connect the results of this work with known mechanisms of the power law [1], such as scale invariance, universality and self-organized criticality.
@@
Acknowledgements.— I would like to thank Tobias Haug, Jessica Jiang, Jingu Pang, John Preskill, Gil Refael and Yu Tong for insightful discussions. I am also particularly grateful to Andy Chia, Leong-Chuan Kwek and Changsuk Noh for collaborations on related projects. The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.
Appendix A Steady state number distribution for the $M$-boson model
@@
[…]
@@
It is straightforward to show that the steady state does not generically obey detailed balance, and is therefore nonequilibrium. Detailed balance imposes the constraint
$\rho_{n,n}\kappa_{m}\prod_{j=1}^{m}(n+j)=\rho_{n+m,n+m}\gamma_{m}\prod_{j=0}^{m-1}(n-j),$ (27)
for all $m=1,\ldots,M$. Substituting the asymptotic behavior (18) for $\rho_{n,n}$, and matching both sides to leading order in $n$, yields the constraint
$\xi=\frac{m}{\ln(\gamma_{m}/\kappa_{m})}.$ (28)
This is only satisfied for $m=M$, and not generally for other values of $m$.
Appendix B Analytical results for $M=2$ at the critical point
Here, we present additional analytical results for the $M$-boson model, in the simplest non-trivial case of $M=2$.
B.1 Moments of the number distribution